OpenStudy (anonymous):

Evaluate the expression and write the result in the form a + bi. (−4 + i)(4 − 5i)

OpenStudy (anonymous):

Foil it out.

OpenStudy (anonymous):

If you see an \(i^2\) replace it with a (-1)

OpenStudy (anonymous):

im not sure. what does it come out too

OpenStudy (lalaly):

-16 +20i +4i+20 = 24i +4

OpenStudy (anonymous):

(-4 + i) (4 - 5 i) = -16 + 20 i + 4 i - 5 (i^2) = -11 + 24 i